OpenStudy (anonymous):

Have to show work and need to write the answer in Scientific Notation 8.4x10^12/-7x10^9

OpenStudy (campbell_st):

well you can do it this way \[\frac{(8.4 \times 10^{10}) \times 10^2}{-7 \times 10^9}\] which is like saying \[\frac{84 \times 10^9}{-7 \times 10^9} \times 10^2\] which leaves \[- \frac{84}{7} \times 10^2\] which is \[-12\times 10^2... or... -1.2 \times 10^3\]

OpenStudy (anonymous):

how did you get 10^2

OpenStudy (campbell_st):

I moved the decimal point 1 place to the right in the number 8.4 to get 84 ... which is the same as multiplying by 10...

OpenStudy (anonymous):

yeah but its 10 to the 12th power